Find the centre of gravity of a thin sheet with density d(x, y) = y, bounded by the

curves y = 4x2 and x = 4.


1
Expert's answer
2021-05-07T09:20:45-0400
"m=\displaystyle\int_{0}^4\displaystyle\int_{0}^{4x^2}ydydx"

"=\displaystyle\int_{0}^4[\dfrac{y^2}{2}]\begin{matrix}
 4x^2 \\
 0
\end{matrix}dx=\displaystyle\int_{0}^48x^4dx=[\dfrac{8x^5}{5}]\begin{matrix}
 4 \\
 0
\end{matrix}"

"=\dfrac{8192}{5}"

"M_x=\displaystyle\int_{0}^4\displaystyle\int_{0}^{4x^2}y^2dydx"

"=\displaystyle\int_{0}^4[\dfrac{y^3}{3}]\begin{matrix}
 4x^2 \\
 0
\end{matrix}dx=\displaystyle\int_{0}^4\dfrac{64x^6}{3}dx=[\dfrac{64x^7}{21}]\begin{matrix}
 4 \\
 0
\end{matrix}"

"=\dfrac{1048576}{21}"

"M_y=\displaystyle\int_{0}^4\displaystyle\int_{0}^{4x^2}xydydx"

"=\displaystyle\int_{0}^4[\dfrac{xy^2}{2}]\begin{matrix}
 4x^2 \\
 0
\end{matrix}dx=\displaystyle\int_{0}^48x^5dx=[\dfrac{8x^6}{6}]\begin{matrix}
 4 \\
 0
\end{matrix}"

"=\dfrac{16384}{3}"


"\bar{x}=\dfrac{M_y}{m}=\dfrac{16384(5)}{3(8192)}=\dfrac{10}{3}"


"\bar{y}=\dfrac{M_x}{m}=\dfrac{1048576(5)}{21(8192)}=\dfrac{640}{21}"



"(\dfrac{10}{3}, \dfrac{640}{21})"